# Region-aware Image-based Human Action Retrieval with Transformers

### v1

@@ Abstract
Human action understanding is a fundamental and challenging task in computer vision. Although there exists tremendous research on this area, most works focus on action recognition, while action retrieval has received less attention. In this paper, we focus on the neglected but important task of image-based action retrieval which aims to find images that depict the same action as a query image. We establish benchmarks for this task and set up important baseline methods for fair comparison. We present an end-to-end model that learns rich action representations from three aspects: the anchored person, contextual regions, and the global image. A novel fusion transformer module is designed to model the relationships among different features and effectively fuses them into an action representation. Experiments on the Stanford-40 and PASCAL VOC 2012 Action datasets show that the proposed method significantly outperforms previous approaches for image-based action retrieval.
@@ 1 Introduction
Understanding human actions from visual data is a crucial task in the computer vision community, with tremendous applications in video surveillance, autonomous driving, human–robot interaction, health monitoring, etc. Action recognition and action retrieval are two important tasks for human action understanding. Recognition aims to classify the action in a given video or image, while retrieval strives to find videos or images that depict the same action as a query video or image.
Despite great progress has been made, most works focus on video-based human action understanding and particularly on the task of video-based action recognition Kong and Fu (2022). Action retrieval has received less attention than action recognition in the academic community. Limited existing works of human action retrieval are primarily targeted at actions in the video Wray et al. (2019); Kico et al. (2022). Although videos contain more information that serves better action recognition, understanding actions in videos requires more computational resources.
@@
Although image-based action retrieval can be seen as a particular application of instance-level image retrieval, it is more challenging than retrieving the landmark or product image. This is due to the fact that understanding human actions involves a semantically higher level of perception than object recognition, and there are often more significant appearance differences among images depicting the same action. Fig. 1 shows the example of image-based action retrieval. Considering the action of ”repairing the bike,” the query image can easily be confused with the action of ”riding the bike,” particularly when the appearances or persons depicted in the image are very similar.
In this paper, we study the neglected but important task of image-based action retrieval. As the image may contain multiple persons, an anchored region is required to indicate the person who performs the action. Since the action involves interactions between the person and contextual objects, we hypothesize that the possible regions located close to the human in the image also play a more important role in understanding human actions.
@@
    To fully exploit person-related and contextual cues for action retrieval, we present an end-to-end model that learns multi-level representations.
@@
    We design a fusion transformer module to fuse different features and obtain an efficient action representation.
  • 4.
@@ 2.1 Image-Based Action Recognition
Image-based action recognition is a challenging task that requires understanding the human as well as its context. Deep learning has been widely applied to this task and achieved remarkable results. We categorize recent works on this topic into three groups: body part-based method, human-object interaction-based method and transfer learning-based method.
@@
Human-object interaction-based method studies the relationship between humans and objects in the scene and use it as a cue for action recognition. For example, Wu and Yu (2021) introduce an efficient relation module that combines human-object and scene-object relations for action recognition. Girish et al. (2020) develop an action recognition method based on target detection and pose estimation that leverages the spatial relationship between humans and objects and the hierarchical relationship between action categories. These methods aim to exploit the contextual information of objects for action recognition.
Transfer learning-based method uses pre-trained models on large-scale datasets and fine-tune them on specific action recognition datasets. For example, Mohammadi et al. (2019) employ the transfer learning technique to address the problem of data scarcity in action recognition datasets. Chakraborty et al. (2021) use a pre-trained model on ImageNet and fine-tune it on a small-scale action recognition dataset using a novel loss function. These methods aim to exploit the generalization ability of pre-trained models for action recognition.
@@ 2.2 Human Action Retrieval
Human action retrieval is a new and challenging topic in computer vision that aims to find videos or images that contain specific human actions. Most existing methods retrieve human actions from videos by describing local features of the motion patterns Veinidis et al. (2019). Kico et al. (2022) present a method based on metric learning to learn a feature space that preserves the similarity of human actions and enables efficient content-based retrieval. Some other methods Veinidis et al. (2019) use unsupervised approaches to cluster and retrieve action videos based on motion features.
@@ 2.3 Instance-level Image Retrieval
Instance-level image retrieval is the task of searching in a large database for images that match an object in a query image. This task requires effective and efficient methods to represent and compare images. Some works Siméoni et al. (2019); Cao et al. (2020) adopt a geometric verification approach to rerank images based on local features. Recently, some methods Radenović et al. (2018); Tan et al. (2021); Wu et al. (2022) have aimed at learning more powerful representations of images using global features.
Siméoni et al. (2019) propose a new Deep Spatial Matching (DSM) method for image retrieval that uses an initial ordering of image descriptors extracted from the activation layer of a convolutional neural network, followed by a set of local features approximating the same sparse 3D activation tensor. Cao et al. (2020) unify global and local features into a single deep model, enabling accurate retrieval with efficient feature extraction. The model combines generalized mean pooling for global features and attentive selection for local features. Radenović et al. (2018) propose a method to fine-tune CNNs for image retrieval on a large collection of unordered images in a fully automated manner. They use hard negative mining and feature normalization to improve the performance of global descriptors. Tan et al. (2021) propose Reranking Transformers as a general model to incorporate both local and global features to rerank the matching images in a supervised fashion and thus replace the relatively expensive process of geometric verification. RRTs are lightweight and can be easily parallelized so that reranking a set of top matching results can be performed in a single forward-pass. Wu et al. (2022) propose a tokenizer module to aggregate deep local features of images into a few visual tokens, then enhance these tokens by a refinement block, and finally generate the global representation by concatenating these visual tokens.
@@ 3 Method
Fig. 2: The overall architecture of the proposed method. Our model is mainly composed of the proposal generation module, the multi-level representations learning module, and the fusion transformer module.
To enhance understanding of human actions within images, we aim to harness representations that span various levels of granularity: the anchored person, contextual regions, and the entire image. The anchored person specifically denotes the human body, whereas contextual regions represent objects proximate to the human that are pertinent to the action being depicted. The overall architecture is depicted in Fig. 2, and the details of the key components are described as follows.
@@ 3.2 Representations of Contextual Regions
where $\mathbb{P}$ denotes the proposals with the confidence scores above the threshold, $\widetilde{\mathbb{P}}$ denotes the selected top-k proposals about contextual regions, $\mathbf{F_{l}}$ is the set of representations of contextual regions.
@@ 3.3 Fusion Transformer
Fig. 3: The architecture of fusion transformer module. Composed of N Transformer blocks.
@@ 4.1 Datasets and Implementation Details
Implementation Details. For proposal generation, we use FasterRCNN Ren et al. (2015) with pre-trained parameters. The backbone of representation learning module is the ResNet-50 pre-trained on ImageNet. The fusion transformer has 3 Transformer blocks with 8 heads each and is trained from scratch. The input size of images is $224\times 224$, with data augmentation including horizontal flipping, random crop, random color jitter, and mix-up during training. We use the batch size of 128 to train our model on a single Nvidia RTX 3090 GPU. The early-stop strategy is used and the max training epochs is set to 150. AdamW is used to optimize the model with an initial learning rate is 0.00003. The number of selected bounding boxes $k$ is set to 10. The feature dimension $D$ is set to 2048.
@@ 4.2 Experimental Setup
Image-based Action Retrieval. As there are no benchmarks for image-based action retrieval, we establish experimental settings and conduct baselines for subsequent analysis and comparison. The experimental settings for image-based action retrieval are constructed based on those of action recognition on the Stanford-40 and PASCAL VOC datasets. Specifically, we train the action representation model on the training set, and consider the validation set as both the query and gallery sets for action retrieval. As the query set and gallery set are the same set, we remove the query sample from the gallery set while retrieving.
@@
Evaluation Metrics. For action retrieval, we adopt the widely used mAP, Rank-1, Rank-5 as the evaluation metrics. Unless otherwise specified, we use the reranking method in Luo et al. (2019), and set the parameters $k1=200$, $k2=20$, and $lambda=0.3$.
Table 1: Results of static action retrieval on the Stanford-40 dataset.
@@ 4.4 Results of Action Recognition
Table 3: Results of image-based action recognition on the Stanford-40.
@@
Ours 89.0
@@ 4.5 Ablation Studies
Fig. 5: Visualization of retrieval results. For each example, the query image is on the left, and the top-$8$ retrieved images are displayed from left to right. The red box denotes the correct retrieved sample. DELG, Baseline, and Token denote DELG global Cao et al. (2020), Strong baseline Luo et al. (2019), Token Wu et al. (2022), respectively, which are representative instance-level image retrieval methods.
@@ 4.6 Visualizations
Fig. 5 shows the retrieval results for different methods. The image on the left is the query, and the red boxes note the correct retrieved samples from the gallery. Our proposed method outperforms other methods in terms of retrieval performance. For the action “repair bicycle,” our method correctly retrieves 7 out of the top 8 results. In a challenging case for the action “Drinking,” our method retrieves only 4 correct samples out of the top 8 results. Our method tends to misclassify some error samples as “smoking” actions, as they share significant similarities with the query action. Nevertheless, in comparison to other methods, the error samples retrieved by our method are semantically more reasonable.
@@ 5 Conclusion
This paper presents a general method for image-based human action understanding, with a special focus on the novel task of image-based action retrieval. We establish two benchmarks for this task and establish important baseline methods for fair comparison. The proposed method learns multi-level features that capture three aspects of an action: the anchored person, contextual regions, and the global image. A novel Transformer-based fusion module is proposed to integrate these features into a enhanced representation. Our approach achieves state-of-the-art performance on both action recognition and retrieval tasks. Furthermore, it demonstrates the capability to retrieve images depicting the same or similar actions as a given query image using a single image. Ablation studies highlight the effectiveness and complementarity of the three types of representations. We hope that this work and the established benchmarks can promote future research on image-based action retrieval.

### v2

@@ Abstract
Human action understanding is a fundamental and challenging task in computer vision. Although there exists tremendous research on this area, most works focus on action recognition, while action retrieval has received less attention. In this paper, we focus on the neglected but important task of image-based action retrieval which aims to find images that depict the same action as a query image. We establish benchmarks for this task and set up important baseline methods for fair comparison. We present a Transformer-based model that learns rich action representations from three aspects: the anchored person, contextual regions, and the global image. A fusion transformer is designed to model the relationships among different features and effectively fuse them into an action representation. Experiments on both the Stanford-40 and PASCAL VOC 2012 Action datasets show that the proposed method significantly outperforms previous approaches for image-based action retrieval.
@@ 1 Introduction
Understanding human actions from visual data is a crucial task in the computer vision community, with tremendous applications in video surveillance, autonomous driving, human–robot interaction, health monitoring, etc. Action recognition and action retrieval are two important tasks for human action understanding. The former aims to classify human action in a given video or image, while the latter strives to find videos or images that depict the same action as a query video or image.
Despite great progress has been made, most works focus on video-based human action understanding and particularly on the task of video-based action recognition Kong and Fu (2022). Action retrieval has received less attention than action recognition in the academic community. Limited existing works of human action retrieval are primarily targeted at actions in the video Wray et al. (2019); Kico et al. (2022a). Although videos contain more information that serves better action recognition, understanding actions in videos requires more computational resources.
@@
Although image-based action retrieval can be seen as a particular application of instance-level image retrieval, it is more challenging than retrieving the landmark or product image. This is due to the fact that understanding human actions involves a semantically higher level of perception than object recognition, and there are often more significant appearance differences among images depicting the same action. Fig. 1 shows the example of image-based action retrieval. Considering the action of “repairing the bike,” the query image can easily be confused with the action of “riding the bike,” particularly when the appearances or persons depicted in the image are very similar. In addition, image-based human action retrieval not only promotes deeper research into visual human action understanding, but also demonstrates significant practical potential in real-time and efficient action recognition/retrieval scenarios.
In this paper, we study the neglected but important task of image-based action retrieval. As the image may contain multiple persons, an anchored region is required to indicate the person who performs the action. Since the action involves interactions between the person and contextual objects, we hypothesize that the possible regions located close to the human in the image also play a more important role in understanding human actions. Considering that the Transformer Vaswani et al. (2017) is capable of learning a joint representation that integrates information from multiple modalities, rendering it highly suitable for fusion tasks, we employ the Transformer to fuse the representations of the person, contextual objects, and their interactions. In the query example of repairing the bike in Fig. 1, simple feature fusion strategies like averaging pooling or concatenation are unable to capture intricate interactions and relationships between the person and related objects. In contrast, Transformer can learn the joint representations that encapsulates the complementary information between them.
@@
    We introduce an efficient Region-aware Image-based human Action Retrieval with Transformers (RIART), which leverages both person-related and contextual object cues, and employs a fusion transformer module for human action retrieval.
@@ 2.1 Image-Based Action Recognition
Image-based action recognition is a complex and challenging endeavor that requires a deep comprehension of the human actions depicted within the image. Deep learning has been widely applied to this task and achieved remarkable results. We categorize recent works on this topic into three groups: body part-based method, human-object interaction-based method and transfer learning-based method.
@@
Human-object interaction-based method studies the relationship between humans and objects in the scene and use it as a cue for action recognition. For example, Wu and Yu (2021) introduce an efficient relation module that integrates both human-object and scene-object relationships for the purpose of enhancing action recognition capabilities. Girish et al. (2020) develop an action recognition method based on target detection and pose estimation that leverages the spatial relationship between humans and objects and the hierarchical relationship between action categories. These methods aim to exploit the contextual information of objects for action recognition.
Transfer learning-based method uses pre-trained models on large-scale datasets and fine-tune them on specific action recognition datasets. For example, Mohammadi et al. (2019) employ the transfer learning technique to address the problem of data scarcity in action recognition datasets. Chakraborty et al. (2021) use a pre-trained model on ImageNet and fine-tune it on a small-scale action recognition dataset using a novel loss function. These methods aim to exploit the generalization ability of pre-trained models for action recognition. Wang et al. (2018) aim to hallucinate temporal characteristics from static images, leveraging video memory techniques, in order to enhance action recognition capabilities with few still images.
To the best of our knowledge, only a very limited number of works Ramanathan et al. (2015); Li et al. (2011) have studied image-based action retrieval, primarily owing to the scarcity of benchmarks. The work in Ramanathan et al. (2015) follows the traditional Bag-of-Words pipeline, which is based on manually designed local feature descriptors. The work Li et al. (2011) aims to learn global image-level action representations by leveraging the semantic relationship between different actions. However, both models are only suitable for simple images that contain a single action and are not scalable enough for large-scale applications. In contrast, we comprehensively study image-based action retrieval and introduce an efficient, deep learning-based pipeline.
@@ 2.2 Human Action Retrieval
Human action retrieval is a new and challenging topic that aims to find videos or images that contain specific human actions. Early works rely on efficient feature matching methods based on local features Tang et al. (2013). Most existing methods retrieve human actions from videos by describing local features of the motion patterns Veinidis et al. (2019). Kico et al. (2022a) present a metric learning-based method to learn a feature space that preserves the similarity of human actions and enables efficient content-based retrieval. Some other methods Veinidis et al. (2019) use unsupervised approaches to cluster and retrieve action videos based on motion features. There are also works Kico et al. (2022b) that focus on content-based action retrieval from human skeleton sequences. It should be noted that human action retrieval is also closely related to zero-shot action recognition C et al. (2021), which aims to recognize novel action categories without training data. However, zero-shot action recognition emphasizes generalization to unseen classes, while action retrieval emphasizes accuracy and efficiency of retrieval.
@@ 2.3 Instance-level Image Retrieval
Instance-level image retrieval involves searching a vast database to retrieve images that contain an object identical to the one specified in a query image. This task requires effective and efficient methods to represent and compare images. Some works Siméoni et al. (2019); Cao et al. (2020) adopt a geometric verification approach to rerank images based on local features. Recently, some methods Radenović et al. (2018); Tan et al. (2021); Wu et al. (2022) have aimed at learning more powerful representations of images using global features.
Siméoni et al. (2019) propose a new Deep Spatial Matching (DSM) method for image retrieval that uses an initial ordering of image descriptors extracted from the activation layer of a convolutional neural network, followed by a set of local features approximating the same sparse 3D activation tensor. Cao et al. (2020) integrate global and local features within a unified deep model, facilitating precise retrieval through efficient feature extraction. This integration is achieved by leveraging generalized mean pooling for capturing holistic global characteristics and attentive selection mechanisms for pinpointing salient local details. Radenović et al. (2018) fine-tune CNNs for image retrieval on extensive collections of unordered images. They use hard negative mining and feature normalization to improve the performance of global descriptors. Tan et al. (2021) introduce Reranking Transformers as a versatile model that incorporates both local and global features to refine the ranking of matched images in a supervised manner, thereby streamlining the typically resource-intensive process of geometric verification. Wu et al. (2022) propose a tokenizer module to aggregate deep local features of images into a few visual tokens, then enhance these tokens by a refinement block, and finally generate the global representation by concatenating these visual tokens.
@@ 3 Method
To enhance understanding of human actions within images, we aim to harness representations that span various levels of granularity: the anchored person, contextual regions, and the entire image. The anchored person specifically denotes the human body, whereas contextual regions represent objects proximate to the human that are pertinent to the action being depicted. We propose an efficient Region-aware Image-based human Action Retrieval with Transformer (RIART). The overall architecture of the proposed method is illustrated in Fig. 2, and the details of its key components are described as follows.
Fig. 2: The overall architecture of the proposed Region-aware Image-based human Action Retrieval with Transformer (RIART). The RIART takes an image and the corresponding anchored region of the human bounding box as input, generating a human action representation that can be utilized for action retrieval or recognition. To acquire contextual regions associated with the action, an off-the-shelf proposal generation module is employed to generate object proposals, which are then filtered and ranked by a contextual regions ranking module. Multi-level action representations are obtained, specifically, the representation of the anchored person $F_{a}$, representations of contextual regions $\mathbf{F_{i}}$, and the global image-level representation $F_{g}$. Subsequently, these representations are fused by a fusion transformer module to yield the final representation $F_{f}$.
3.1 Problem Formulation
Fig. 3: Examples query image for action retrieval. Human bounding boxes indicate the person performing a certain action. (a) Simple image that contains only one human action. (b) Complex image that contains multiple actions from different persons. Note that actions of different persons in the same image can be different.
Image-based human action retrieval refers to the process of retrieving images from a database that depict specific human actions. Unlike video-based human action retrieval, the actions of interest for this task are typically static actions that can be exemplified in a single image. Examples of query image are shown in Fig. 3. The query image may contain multiple action classes within the same image. We assume that a given human bounding box in an image only encompasses one action class. If the query image depicts multiple persons performing different actions, each image with distinct human bounding boxes is regarded as a separate query example.
@@ 3.3 Representations of Contextual Regions
where $\mathbb{P}$ denotes the proposals with the confidence scores above the threshold, $\widetilde{\mathbb{P}}$ denotes the selected top-k proposals about contextual regions, $\mathbf{F_{l}}$ is the set of representations of contextual regions (for symbols’ fonts, italic and bold styles are used to denote feature vectors and sets of features, respectively).
@@ 3.4 Fusion Transformer
It should be noted that Transformer-based fusion has also appeared in recent works Liu et al. (2022) for multi-modal human action recognition. For instance, a cross-modality fusion module is applied to learn better multi-modal representations Liu et al. (2022). Unlike the fusion modules of these approaches, the proposed Fusion Transformer aims to learn efficient representations from a single image modality by fusing multi-level or multi-grained features. Additionally, type embedding and position embedding are designed to facilitate the feature learning.
Fig. 4: The architecture of Fusion Transformer module. The input $F_{c}$ denotes multi-level action representations, i.e., representation of the anchored person $F_{a}$, representations of contextual regions $\mathbf{F_{i}}$, and the global image-level representation $F_{g}$. The Fusion Transformer aims to capture the complementary information among these diverse input representations and output a learned joint representation. To facilitate the learning of this complementary action representation, type embeddings and position embeddings are employed. Similar to the standard Transformer architecture, the Fusion Transformer module consists of $N$ blocks.
@@ 4.1 Datasets and Implementation Details
Implementation Details. For proposal generation, we use FasterRCNN Ren et al. (2015) with pre-trained parameters. The default backbone for the representation learning module is ResNet-50 He et al. (2016). Additionally, a Transformer-based backbone, specifically Swin-T Liu et al. (2021), is also employed due to its comparable model size to ResNet-50. The fusion transformer has 3 Transformer blocks with 8 heads each and is trained from scratch. The input size of images is $224\times 224$, with data augmentation including horizontal flipping, random crop, random color jitter, and mix-up during training. We use the batch size of 128 to train our model on a single Nvidia RTX 3090 GPU. The early-stop strategy is used and the max training epochs is set to 150. AdamW is used to optimize the model with an initial learning rate is 0.00003. The number of selected bounding boxes $k$ is set to 10. The feature dimension $D$ is set to 2048.
@@ 4.2 Experimental Setup
Image-based Action Retrieval. As there are no benchmarks for image-based action retrieval, we establish experimental settings and conduct baselines for subsequent analysis and comparison. The experimental settings for image-based action retrieval are constructed based on those of action recognition on both the Stanford-40 and PASCAL VOC datasets. Specifically, we train the action representation model on the training set, and consider the validation set as both the query and gallery sets for action retrieval. As the query set and gallery set are the same set, we remove the query sample from the gallery set while retrieving. The query set refers to the collection of query images that are used as input to the image-based retrieval model, while the gallery set is a collection of images that the image retrieval system searches through to find matches for the query images.
@@
Evaluation Metrics. For action retrieval, we adopt the widely used mAP, Rank-1, Rank-5 as the evaluation metrics. Reranking Luo et al. (2019); Zhang et al. (2024) is a common post-processing step of reordering the initial retrieved results based on additional criteria, with the goal of improving the retrieved results. Unless otherwise specified, we use the reranking method in Luo et al. (2019), and set the parameters $k1=200$, $k2=20$, and $lambda=0.3$.
Table 1: Results of human action retrieval on the Stanford-40 dataset. In order to keep the model lightweight, we employ not only ResNet50 but also Swin-T Liu et al. (2021) as the backbone. The choice of Swin-T stems from the fact that its model size is comparable to that of ResNet50.
@@ 4.4 Results of Action Recognition
Table 3: Results of image-based action recognition on the Stanford-40. The symbol “*” signifies human pose-based methods that utilize human pose as auxiliary information, whereas the alternative approach solely necessitates the global human bounding box.
@@
MASPP Ashrafi et al. (2023)* 94.8
RIART (ResNet-50) 89.0
RIART (Swin-T) 91.8
@@ 4.5 Ablation Studies
Fig. 6: Visualization of retrieval results. For each example, the query image is on the left, and the top-$8$ retrieved images are displayed from left to right. The red box indicates the correctly retrieved sample, whereas the other images that do not have red boxes represent wrongly retrieved samples. DELG, Baseline, and Token denote DELG global Cao et al. (2020), Strong baseline Luo et al. (2019), Token Wu et al. (2022), respectively, which are representative instance-level image retrieval methods. The anchored human box in the query image is omitted for simplicity. The first four, i.e., “blowing bubbles”, “drinking”, “fixing a bike” and “reading”, and our retrieved results are mostly accurate. The middle two, i.e., “pouring liquid” and “texting message” pose a bit of a challenge, and our approach can only correctly retrieve two out of the top-8 results. The last four, i.e., “playing guitar”, “writing on a book”, “fishing” and “phoning”, are hard examples, and most approaches almost fail.
@@ 4.6 Visualizations
Fig. 6 shows the retrieval results for different methods. The image on the left is the query, and the red boxes note the correct retrieved samples from the gallery. Our proposed method outperforms other methods in terms of retrieval performance. For the action “repair bicycle”, our method correctly retrieves 7 out of the top-8 results. In a challenging case for the action “Drinking”, our method retrieves only 4 correct samples out of the top-8 results. Our method tends to misclassify some error samples as “smoking” actions, as they share significant similarities with the query action. The last four, i.e., “playing guitar”, “writing on a book”, “fishing” and “phoning”, represent challenging examples for image-based action retrieval, and most approaches only achieve unsuccessful retrieval results. Nevertheless, in comparison to other methods, the error samples retrieved by our method are semantically more reasonable.
@@ 5 Conclusion
In this paper, we empirically investigate the overlooked task of image-based human action retrieval, and establish new benchmarks and important baselines for this task. We introduce an efficient Region-aware Image-based human Action Retrieval with Transformers (RIART), which learns multi-level action representations from three perspectives: the anchored person, contextual regions, and the global image. A Transformer-based fusion module is also proposed to integrate these features into an enhanced representation. Our approach achieves state-of-the-art performance on both action recognition and retrieval tasks. Furthermore, it demonstrates the capability to retrieve images depicting the same or similar actions as a given query image using a single image. Ablation studies highlight the effectiveness and complementarity of the three types of representations. We hope that this work and the established benchmarks can promote future research on image-based action retrieval.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (62302093, 62172090, 62202438), Jiangsu Province Natural Science Fund (BK20230833), CAAI-Huawei MindSpore Open Fund, and the Southeast University Start-Up Grant for New Faculty (RF1028623063, RF1028623097). This work is also supported by the Big Data Computing Center of Southeast University.